9. A ladder 20 ft long is leaning against an embankment inclined 60 ° to the horizontal. If the bottom of the ladder is moved horizontally toward the embankment at 1 ft/sec, how fast is the top of the ladder moving when the bottom is 4 ft from the embankment?
1
Expert's answer
2021-10-17T17:45:15-0400

Consider the diagram below


Let AC be x ft and BC be y ft

"x^2+y^2=20^2\\
y^2=20^2-x^2\\
2y \frac{dy}{dt}=-2x\frac{dx}{dt}\\
Also \\
\frac{dx}{dt}=1 \implies 2y \frac{dy}{dt}=-2x*1 \implies \frac{dy}{dt}= -\frac{x}{y}\\
y^2=400-16 \implies y =\sqrt{384}\\
\frac{dy}{dt}= -\frac{4}{\sqrt{384}}=-0.2041"

Therefore the top of the ladder is moving downwards at 0.2041 ft/s